# Small-Time Local Controllability of the multi-input bilinear Schrödinger equation thanks to a quadratic term

### v3

@@ Abstract
We focus on the small-time local controllability (STLC) around an equilibrium, when the linearized system is not controllable. We study the second order term in the Taylor expansion of the state, with respect to the control.
@@ 1.1 Model and problem
This equation describes the evolution of the wave function $\psi$ of a quantum particle, in an $1$D infinite square potential well $(0,1)$, subjected to two electric fields with amplitudes $u(t)$ and $v(t)$. The functions $\mu_{1}$ and $\mu_{2}$, called "dipolar moments", model the interaction between the particle’s wave function $\psi$ and the two electric fields $u,v$.
This is multi-input nonlinear control system:
@@
    the control is $(u,v):(0,T)\to\mathbb{R}^{2}$, it acts bilinearly on the state, through the term $(u(t)\mu_{1}(x)+v(t)\mu_{2}(x))\psi(t,x)$, this is a reason why we speak of "bilinear Schrödinger equation".
The ground state is the particular trajectory $\psi_{1}(t,x):=\varphi_{1}(x)e^{-i\lambda_{1}t}$, with $(u,v)=0$ where $\varphi_{1}(x):=\sqrt{2}\sin(\pi x)$ and $\lambda_{1}:=\pi^{2}$. We are interested in the local controllability of the nonlinear system (1.1), around the ground state, i.e. realizing small motions around the ground state with small controls: given a maximal size $\varepsilon>0$ for the controls, does there exist $T>0$ and $\delta>0$ such that, for any target $\psi_{f}$ close enough to the ground state, i.e. $\|\psi_{f}-\psi_{1}(T)\|<\delta$, there exist controls $(u,v):(0,T)\to\mathbb{R}^{2}$ such that $\|(u,v)\|<\varepsilon$ and $\psi(T;(u,v),\varphi_{1})=\psi_{f}$. The different norms used are of great importance and will be precised later.
More precisely, we are interested in the small-time local controllability around the ground state i.e. realizing, in arbitrary small times $T>0$, small motions around the ground state with small controls: given $T,\varepsilon>0$, does there exist $\delta>0$ such that, for any target $\psi_{f}$ such that $\|\psi_{f}-\psi_{1}(T)\|<\delta$, there exist controls $(u,v):(0,T)\to\mathbb{R}^{2}$ such that $\|(u,v)\|<\varepsilon$ and $\psi(T;(u,v),\varphi_{1})=\psi_{f}$. It corresponds to the local surjectivity around 00 of the nonlinear end-point map $\Theta_{T}:(u,v)\mapsto\psi(T;(u,v),\varphi_{1})$.
@@ 1.2 The controllable linearized case: constrained functional framework
Classical functional analysis study shows that the eigenvalues and eigenvectors are
@@
$\psi_{j}(t,x):=\varphi_{j}(x)e^{-i\lambda_{j}t},\qquad(t,x)\in(0,T)\times(0,1),\qquad j\geqslant 1.$
These functions are the solutions to the free (with controls $(u,v)\equiv 0$) Schrödinger equation (1.1), with initial data $\varphi_{j}$ at time $t=0$. They are called eigenstates. When $j=1$, $\psi_{1}$ is called the fundamental state, or ground state. We finally define the spaces $H^{s}_{(0)}(0,1)=D(A^{s/2})$, equipped with the norm
@@
Let $p,m\in\mathbb{N}$ and $\mu_{1},\mu_{2}\in H^{2(p+m)+3}(0,1)$ be such that $\mu_{\ell}^{(2k+1)}(0)=\mu_{\ell}^{(2k+1)}(1)=0$ for $k=0,\cdots,p-1$, $\ell=1,2$ and
@@
Then the Schrödinger equation (1.1) is $H^{m}_{0}((0,T),\mathbb{R})-$small-time locally controllable in $H^{2(p+m)+3}_{(0)}(0,1)$: for every $T,\varepsilon>0$, there exists $\delta>0$ such that for every $\psi_{f}\in\mathcal{S}\cap H^{2(p+m)+3}_{(0)}(0,1)$ with $\|\psi_{f}-\psi_{1}(T)\|_{H^{2(p+m)+3}}<\delta$, there exists $(u,v)\in H^{m}_{0}((0,T),\mathbb{R})^{2}$ such that $\|(u,v)\|_{H^{m}}\leqslant\varepsilon$ and $\psi(T;(u,v),\varphi_{1})=\psi_{f}$.
@@
is necessary for the controllability of the linearized system around the ground state. The stronger assumption (1.4) guarantees that the linearized system is $H^{m}_{0}((0,T),\mathbb{R})$-small-time controllable in $H^{2(p+m)+3}(0,1)$. Then, a smoothing effect allows to prove that the end-point map is of class $\mathcal{C}^{1}$ between the following spaces
@@
Note that the assumption (1.4) prevents the $\mu_{\ell}^{(2p+1)}$ from vanishing at the boundary because, if $\mu_{\ell}$ satisfies the hypotheses of Theorem 1.1, then integration by parts give: for all $k,j\in\mathbb{N}^{*}$, for all $\ell\in\{1,2\}$,
@@ 1.3 Our case study and its interpretation in terms of Lie brackets
In this case, the linearized system is not controllable: it misses one complex direction: $\langle\psi(t),\varphi_{K}\rangle\in\mathbb{C}$, i.e. the first order term in the Taylor expansion at $(0,0)$ of the map $(u,v)\mapsto\langle\psi(t;(u,v),\varphi_{1}),\varphi_{K}\rangle$ vanishes. We assume that this is the only direction lost, i.e. all other directions are controllable on the linearized system, with the usual functional framework, so we consider integers $p,m\in\mathbb{N}$ and we assume
@@
To control the complex direction $\langle\psi(t),\varphi_{K}\rangle$ of the nonlinear system (1.1), we will use a power series expansion of the order $2$, i.e. the second order term in the Taylor expansion at $(0,0)$ of the map $(u,v)\mapsto\langle\psi(t;(u,v),\varphi_{1}),\varphi_{K}\rangle$. At this step of the proof, we consider an integer $n\in\mathbb{N}^{*}$ and we assume that the following assumptions hold:
@@
$A_{k}^{1}:=(-1)^{k-1}\sum_{j=1}^{+\infty}\left(\lambda_{j}-\frac{\lambda_{1}+\lambda_{K}}{2}\right)(\lambda_{K}-\lambda_{j})^{k-1}(\lambda_{j}-\lambda_{1})^{k-1}c_{j}=0.$
@@
$A_{k}^{2}:=(-1)^{k-1}\sum_{j=1}^{+\infty}\left(\lambda_{j}-\frac{\lambda_{1}+\lambda_{K}}{2}\right)(\lambda_{K}-\lambda_{j})^{k-1}(\lambda_{j}-\lambda_{1})^{k-1}\tilde{c_{j}}=0.$
@@
$\gamma_{k}:=\displaystyle\sum_{j=1}^{+\infty}\left((\lambda_{K}-\lambda_{j})^{\lfloor\frac{k+1}{2}\rfloor}(\lambda_{j}-\lambda_{1})^{\lfloor\frac{k}{2}\rfloor}d_{j}-(\lambda_{K}-\lambda_{j})^{\lfloor\frac{k}{2}\rfloor}(\lambda_{j}-\lambda_{1})^{\lfloor\frac{k+1}{2}\rfloor}\tilde{d_{j}}\right)=0.$
@@
For $A$ and $B$ two operators, we define by induction on $k\in\mathbb{N}$ the operator $\underline{\operatorname{ad}}^{k}_{A}(B)$ as: $\underline{\operatorname{ad}}_{A}^{0}(B)=B$ and $\underline{\operatorname{ad}}_{A}^{k+1}(B)=\underline{\operatorname{ad}}^{k}_{A}(B)A-A\underline{\operatorname{ad}}^{k}_{A}(B)$. Under appropriate relations between the parameter $n$ and $p$, the assumptions $\mathbf{(H)_{quad,K,1}}$, $\mathbf{(H)_{quad,K,2}}$, $\mathbf{(H)_{quad,K,3}}$ and $\mathbf{(H)_{quad,K,4}}$ can be interpreted in terms of Lie brackets:
@@
where $A$ is defined in (1.2), and $B_{\ell}$ is the multiplication operator by $\mu_{\ell}$ in $L^{2}(0,1)$. We refer to Propositions A.6, A.7, A.9 and A.10 in appendix for a precise proof.
@@
$u_{0}=u,\qquad\text{and}\qquad\forall n\in\mathbb{N},\ \forall t\in[0,T],\quad u_{n+1}(t)=\int_{0}^{T}u_{n}(t)\mathrm{d}t.$
Then, heuristically, it is as if the following terms were included in the quadratic development of the solution:
@@
The cancellation assumptions $\mathbf{(H)_{quad,K,1}}$, $\mathbf{(H)_{quad,K,2}}$, $\mathbf{(H)_{quad,K,3}}$ reduce this sum to the simpler expression:
@@
and they allow use to take advantage of the term $\int_{0}^{T}u_{\lfloor\frac{n}{2}\rfloor+1}(t)v_{\lfloor\frac{n+1}{2}\rfloor}(t)\mathrm{d}t$. For this strategy to work, it is thus necessary to consider $K\geqslant 2$. Indeed, the assumption $\mathbf{(H)_{lin,K=1,2}}$ leads to $\max(A_{k}^{1},A_{k}^{2})>0$. Thus, at least one of the assumptions $\mathbf{(H)_{quad,1,1}},\mathbf{(H)_{quad,1,2}}$ is not verified.
Therefore, contrary to the single-input case [18], here, in the multi-input case, the quadratic term alone can recover small-time local controllability.
@@ 1.4 Main result and proof strategy
Our main result is the following one, i.e. the same conclusion as in Theorem 1.1, when replacing the assumption (1.4) related to the controllability of the linearized system, by the assumptions $\mathbf{(H)_{lin,K,2}}$,$\mathbf{(H)_{quad,K,1}}$,$\mathbf{(H)_{quad,K,2}}$,$\mathbf{(H)_{quad,K,3}}$, $\mathbf{(H)_{quad,K,4}}$ related to a power series expansion of order two.
@@
Let $p,m,K\in\mathbb{N}$, $n\in\mathbb{N}^{*}$ with $K\geqslant 2$ and $\lfloor\frac{n}{2}\rfloor\leqslant p$, $\mu_{1},\mu_{2}$ satisfying $\mathbf{(H)_{reg}}$, $\mathbf{(H)_{lin,K,1}}$, $\mathbf{(H)_{lin,K,2}}$, $\mathbf{(H)_{quad,K,1}}$, $\mathbf{(H)_{quad,K,2}}$, $\mathbf{(H)_{quad,K,3}}$, $\mathbf{(H)_{quad,K,4}}$ and $\langle\mu_{1}\varphi_{1},\varphi_{1}\rangle\neq 0$. Then the bilinear Schrödinger equation (1.1) is $H^{m}_{0}((0,T),\mathbb{R})-$small-time locally controllable in $H^{2(p+m)+3}_{(0)}(0,1)$.
@@
and the orthogonal projection on $\mathcal{H}$, which is given by
@@
Then, the proof of Theorem 1.3 is divided in two steps: given a time $T>0$, a target $\psi_{f}$ and $z\in\mathbb{C}$,
@@
    we find $T_{1}<T$ and controls $(u,v)$ such that $\langle\psi(T_{1};(u,v),\varphi_{1}),\psi_{K}(T_{1})\rangle\sim z$,
@@
    we apply on $(T_{1},T)$ a control such that $\mathbb{P}_{\mathcal{H}}\psi(T)=\mathbb{P}_{\mathcal{H}}\psi_{f}$,
  3. -
    we prove that, despite the second step, $\langle\psi(T;(u,v),\varphi_{1}),\psi_{K}(T)\rangle$ stays close to $z$, so that a fixed point argument allows to find $z=z(\psi_{f})$ so that $\langle\psi(T;(u,v),\varphi_{1}),\psi_{K}(T)\rangle=\langle\psi_{f},\psi_{K}(T)\rangle.$
Finally $\psi(T;(u,v),\varphi_{1})=\psi_{f}$.
For the step $1$, we examine the quadratic term of $(u,v)\mapsto\langle\psi(t;(u,v),\varphi_{1}),\psi_{K}(t)\rangle$: it contains terms of the form (1.11).

### v4

@@ Abstract
We focus on the small-time local controllability (STLC) around an equilibrium, when the linearized system is not controllable. We study the second-order term in the Taylor expansion of the state, with respect to the control.
@@ 1.1 Model and problem
This equation describes the evolution of the wave function $\psi$ of a quantum particle in a $1$D infinite square potential well $(0,1)$, subjected to two electric fields with amplitudes $u(t)$ and $v(t)$. The functions $\mu_{1}$ and $\mu_{2}$, called "dipolar moments", model the interaction between the particle’s wave function $\psi$ and the two electric fields $u,v$.
This is a multi-input nonlinear control system:
@@
    the controls are $u,v:(0,T)\to\mathbb{R}$, they act bilinearly on the state, through the term $(u(t)\mu_{1}(x)+v(t)\mu_{2}(x))\psi(t,x)$, this is a reason why we speak of "bilinear Schrödinger equation".
The ground state is the particular trajectory $\psi_{1}(t,x):=\varphi_{1}(x)e^{-i\lambda_{1}t}$, with $(u,v)=0$ where $\varphi_{1}(x):=\sqrt{2}\sin(\pi x)$ and $\lambda_{1}:=\pi^{2}$. We are interested in the local controllability of the nonlinear system (1.1) around the ground state, i.e. realizing small motions around the ground state with small controls: given a maximal size $\varepsilon>0$ for the controls, does there exist $T,\delta>0$ such that, for any target $\psi_{f}$ close enough to the ground state, i.e. $\|\psi_{f}-\psi_{1}(T)\|<\delta$, there exist controls $u,v:(0,T)\to\mathbb{R}$ such that $\|(u,v)\|\leqslant\varepsilon$ and $\psi(T;(u,v),\varphi_{1})=\psi_{f}$. The different norms used are of great importance and will be precised later.
More precisely, we are interested in the small-time local controllability around the ground state i.e. realizing, in arbitrarily small times $T>0$, small motions around the ground state with small controls: given $T,\varepsilon>0$, does there exist $\delta>0$ such that, for any target $\psi_{f}$ such that $\|\psi_{f}-\psi_{1}(T)\|<\delta$, there exist controls $u,v:(0,T)\to\mathbb{R}$ such that $\|(u,v)\|\leqslant\varepsilon$ and $\psi(T;(u,v),\varphi_{1})=\psi_{f}$. This corresponds to the local surjectivity around 00 of the nonlinear end-point map $\Theta_{T}:(u,v)\mapsto\psi(T;(u,v),\varphi_{1})$.
@@ 1.2 The controllable linearized case: constrained functional framework
According to classical functional analysis, the eigenvalues and eigenvectors are
@@
$\psi_{j}(t,x):=\varphi_{j}(x)e^{-i\lambda_{j}t},\qquad(t,x)\in(0,T)\times(0,1),\qquad j\geqslant 1.$ (1.4)
These functions are the solutions to the free (with controls $(u,v)\equiv 0$) Schrödinger equation (1.1), with initial data $\varphi_{j}$ at time $t=0$. They are called eigenstates. When $j=1$, $\psi_{1}$ is called the fundamental state, or ground state. Finally, we define the spaces $H^{s}_{(0)}(0,1)=D(A^{s/2})$, equipped with the norm
@@
Let $p,m\in\mathbb{N}$ and $\mu_{1},\mu_{2}\in H^{2(p+m)+3}((0,1),\mathbb{R})$ be such that $\mu_{\ell}^{(2k+1)}(0)=\mu_{\ell}^{(2k+1)}(1)=0$ for $k=0,\cdots,p-1$, $\ell\in\{1,2\}$ and
@@
Then the Schrödinger equation (1.1) is $H^{m}_{0}((0,T),\mathbb{R})-$small-time locally controllable in $H^{2(p+m)+3}_{(0)}(0,1)$: for every $T,\varepsilon>0$, there exists $\delta>0$ such that for every $\psi_{f}\in\mathcal{S}\cap H^{2(p+m)+3}_{(0)}(0,1)$ with $\|\psi_{f}-\psi_{1}(T)\|_{H^{2(p+m)+3}}<\delta$, there exist $u,v\in H^{m}_{0}((0,T),\mathbb{R})$ such that $\|(u,v)\|_{H^{m}}\leqslant\varepsilon$ and $\psi(T;(u,v),\varphi_{1})=\psi_{f}$.
@@
is necessary for the controllability of the linearized system around the ground state. The stronger assumption (1.5) guarantees that the linearized system is $H^{m}_{0}((0,T),\mathbb{R})$-small-time controllable in $H^{2(p+m)+3}_{(0)}(0,1)$. Then, a smoothing effect allows to prove that the end-point map is of class $\mathcal{C}^{1}$ between the following spaces
@@
Note that the assumption (1.5) prevents $\mu_{\ell}^{(2p+1)}$ from vanishing at the boundary because, if $\mu_{\ell}$ satisfies the hypotheses of Theorem 1.1, then integration by parts give: for all $k,j\in\mathbb{N}^{*}$, distinct, for all $\ell\in\{1,2\}$,
@@ 1.3 Our case study and its interpretation in terms of Lie brackets
In this case, the linearized system is not controllable: it misses one complex direction $\langle\psi(t),\varphi_{K}\rangle\in\mathbb{C}$. In other words, the first-order term in the Taylor expansion at $(0,0)$ of the map $(u,v)\mapsto\langle\psi(t;(u,v),\varphi_{1}),\varphi_{K}\rangle$ vanishes. We assume that this is the only direction lost, i.e. all other directions are controllable on the linearized system, with the usual functional framework, so we consider integers $p,m\in\mathbb{N}$ and we assume
@@
To control the complex direction $\langle\psi(t),\varphi_{K}\rangle$ of the nonlinear system (1.1), we will use a power series expansion of the order $2$, i.e. the second-order term in the Taylor expansion at $(0,0)$ of the map $(u,v)\mapsto\langle\psi(t;(u,v),\varphi_{1}),\varphi_{K}\rangle$. At this step of the proof, we consider an integer $n\in\mathbb{N}^{*}$ and we assume that the following assumptions hold:
@@
$A_{k}^{1}:=(-1)^{k-1}\sum_{j=1}^{+\infty}\left(\lambda_{j}-\frac{\lambda_{1}+\lambda_{K}}{2}\right)(\lambda_{K}-\lambda_{j})^{k-1}(\lambda_{j}-\lambda_{1})^{k-1}c_{j}=0,$
@@
$A_{k}^{2}:=(-1)^{k-1}\sum_{j=1}^{+\infty}\left(\lambda_{j}-\frac{\lambda_{1}+\lambda_{K}}{2}\right)(\lambda_{K}-\lambda_{j})^{k-1}(\lambda_{j}-\lambda_{1})^{k-1}\tilde{c_{j}}=0,$
@@
$\gamma_{k}:=\displaystyle\sum_{j=1}^{+\infty}\left((\lambda_{K}-\lambda_{j})^{\lfloor\frac{k+1}{2}\rfloor}(\lambda_{j}-\lambda_{1})^{\lfloor\frac{k}{2}\rfloor}d_{j}-(\lambda_{K}-\lambda_{j})^{\lfloor\frac{k}{2}\rfloor}(\lambda_{j}-\lambda_{1})^{\lfloor\frac{k+1}{2}\rfloor}\tilde{d_{j}}\right)=0,$
@@
For $A$ and $B$ two operators, we define by induction on $k\in\mathbb{N}$ the operator $\underline{\operatorname{ad}}^{k}_{A}(B)$ as: $\underline{\operatorname{ad}}_{A}^{0}(B)=B$ and $\underline{\operatorname{ad}}_{A}^{k+1}(B)=\underline{\operatorname{ad}}^{k}_{A}(B)A-A\underline{\operatorname{ad}}^{k}_{A}(B)$. Under appropriate relations between the parameter $n$ and $p$, the assumptions $\mathbf{(H)_{quad,K,1}}$, $\mathbf{(H)_{quad,K,2}}$, $\mathbf{(H)_{quad,K,3}}$ and $\mathbf{(H)_{quad,K,4}}$ can be interpreted in terms of Lie brackets. More precisely,
@@
where $A$ is defined in (1.2), and $B_{\ell}$ is the multiplication operator by $\mu_{\ell}$ in $L^{2}(0,1)$. We refer to Propositions A.7, A.8, A.10 and A.11 in appendix for a precise proof.
@@
$u_{0}=u,\qquad\text{and}\qquad\forall n\in\mathbb{N},\ \forall t\in[0,T],\quad u_{n+1}(t)=\int_{0}^{t}u_{n}(s)\mathrm{d}s.$
Then, heuristically, it is as if the following terms were included in the quadratic development of the solution
@@
The cancellation assumptions $\mathbf{(H)_{quad,K,1}}$, $\mathbf{(H)_{quad,K,2}}$, $\mathbf{(H)_{quad,K,3}}$ reduce this sum to the simpler expression
@@
and they allow use to take advantage of the term $\int_{0}^{T}u_{\lfloor\frac{n}{2}\rfloor+1}(t)v_{\lfloor\frac{n+1}{2}\rfloor}(t)\mathrm{d}t$. For this strategy to work, it is thus necessary to consider $K\geqslant 2$. Indeed, the assumption $\mathbf{(H)_{lin,1,2}}$ leads to $\max(A_{k}^{1},A_{k}^{2})>0$ for every $k$. Thus, at least one of the assumptions $\mathbf{(H)_{quad,1,1}},\mathbf{(H)_{quad,1,2}}$ is not verified.
Unlike the single-input case [18], the quadratic term alone can recover small-time local controllability in the multi-input case.
@@ 1.4 Main result and proof strategy
Our main result is the following one, i.e. the same conclusion as in Theorem 1.1, when replacing the assumption (1.5) related to the controllability of the linearized system, by the assumptions $\mathbf{(H)_{lin,K,2}}$,$\mathbf{(H)_{quad,K,1}}$,$\mathbf{(H)_{quad,K,2}}$,$\mathbf{(H)_{quad,K,3}}$, and $\mathbf{(H)_{quad,K,4}}$ related to a power series expansion of order two.
@@
Let $p,m,K\in\mathbb{N}$, $n\in\mathbb{N}^{*}$ with $K\geqslant 2$ and $\lfloor\frac{n}{2}\rfloor\leqslant p$. Assume that $\mu_{1},\mu_{2}$ satisfy $\mathbf{(H)_{reg}}$, $\mathbf{(H)_{lin,K,1}}$, $\mathbf{(H)_{lin,K,2}}$, $\mathbf{(H)_{quad,K,1}}$, $\mathbf{(H)_{quad,K,2}}$, $\mathbf{(H)_{quad,K,3}}$, $\mathbf{(H)_{quad,K,4}}$ and $\langle\mu_{1}\varphi_{1},\varphi_{1}\rangle\neq 0$. Then the bilinear Schrödinger equation (1.1) is $H^{m}_{0}((0,T),\mathbb{R})-$small-time locally controllable in $H^{2(p+m)+3}_{(0)}(0,1)$.
@@
and $\mathbb{P}_{\mathcal{H}}$ is the orthogonal projection on $\mathcal{H}$, which is given by
@@
Then, the proof of Theorem 1.3 is divided in two stages: given a time $T>0$, a target $\psi_{f}$ and $z\in\mathbb{C}$,
@@
    we find $T_{1}<T$ and controls $(u_{z},v_{z})$ such that $\langle\psi(T_{1};(u_{z},v_{z}),\varphi_{1}),\psi_{K}(T_{1})\rangle$ is close to $z$,
@@
    we apply on $(T_{1},T)$ controls $(\tilde{u}_{z},\tilde{v}_{z})$ in projection such that on has
    $\mathbb{P}_{\mathcal{H}}\psi(T;(\tilde{u}_{z},\tilde{v}_{z}),\psi(T_{1};(u_{z},v_{z}),\varphi_{1}))=\mathbb{P}_{\mathcal{H}}\psi_{f}.$
    Then, we prove that, despite the second step, $\langle\psi(T;(\tilde{u}_{z},\tilde{v}_{z}),\psi(T_{1};(u_{z},v_{z}),\varphi_{1})),\psi_{K}(T)\rangle$ stays close to $z$, so that a fixed point argument allows to find $z=z(\psi_{f})$ such that $\langle\psi(T;(\tilde{u},\tilde{v}),\psi(T_{1};(u,v),\varphi_{1})),\psi_{K}(T)\rangle=\langle\psi_{f},\psi_{K}(T)\rangle.$
Finally $\psi(T;(\tilde{u},\tilde{v}),\psi(T_{1};(u,v),\varphi_{1}))=\psi_{f}$. We recall that $\psi_{K}$ is defined in (1.4).
In step $1$, we examine the quadratic expansion of $(u,v)\mapsto\langle\psi(T;(u,v),\varphi_{1}),\psi_{K}(t)\rangle$. It contains terms of the form (1.12).